OpenStudy (anonymous):

3b-2=7b+4 what I have is..please tell me if im correct. 3b-2=7b+4 +2 3b=7b+6 -7b -4b=6 1-4 b=(6/-4)=31-2

OpenStudy (anonymous):

yes that´s correct

OpenStudy (tkhunny):

First, what does "31-2" mean? Very odd-looking. Second, what does "1-4" mean. Very odd-looking. Third, and this will be hard. There is no actual reason for the "b" to end up on the left. b = 3 is no better or worse than 3 = b. You can use this to your advantage and make your life a little simpler. 3b-2=7b+4 Subtract 3b -2 = 4b + 4 Subtract 4 -6 = 4b Divide by 2 (Just to simplify) -3 = 2b Divide by 2 -3/2 = b You appear to have the correct answer, but it is very difficult to see how you gto there. Try to be more clear and systematic.

OpenStudy (mathlegend):

I agree with @tkhunny

OpenStudy (tkhunny):

It's probably also important to point out that there is only a visual difference between these expressions: \(-\dfrac{3}{2} = \dfrac{-3}{2} = \dfrac{3}{-2} = -\dfrac{-3}{-2}\)

OpenStudy (anonymous):

3b-2=7b+4 (group by like terms) 3b-7b=4+2 -4b=6 (divide both sides by -4b to get b on its own) b=-6/4 (you can reduce this so you get..0 b=-3/2

OpenStudy (tkhunny):

The in-line notation needs a little help. -3/-2 is troubling. Clarify with parentheses. -3/(-2). It's kind of cute that the leading negative sign is ambiguous. Is it in front of the whole thing or just the numerator? It doesn't matter!

OpenStudy (anonymous):

thank you all for your help